A 45 kg wagon is at rest on a rough surface. The coefficient of sliding friction between the wagon and the floor is 0.80. How much force is needed to start the wagon ?.



1
Expert's answer
2021-04-12T07:05:03-0400

The force, needed to start the wagon should be larger than the static force of friction, which is "F_f = \mu N", where "N" is the normal force, and "\mu" is the coefficient of friction. Here, the normal force is equal to force of gravity, i.e. "N = mg". Hence, "F = 0.8 \cdot 45 kg \cdot 9.81 \frac{m}{s^2} \approx 353.2 N".